There are 174 words typed in 3 minutes. There were 348 words typed in 6 minutes. How many words will be typed in 8 minutes?
Answer by checkley77(12844)

174/3=58 words per minute.
348/6=58 ditto
8*58=464 words can be typed in 8 minutes.